# Maximum Dispersion, Maximum Concentration: Enhancing the Quality of MOP Solutions

### v2

@@ 2 Background
$PF$ =$=\{\mathbf{y}\in F(\Omega);\nexists~{}\mathbf{x}\in\Omega\text{ such that }F(\mathbf{x})\prec\mathbf{y}\}$ (2)
$PO$ =$=\{\mathbf{x}\in\Omega;F(\mathbf{x})\in PF\}$ (3)
The set $PF\subset F(\Omega)$ is called the Pareto Front, and $PO\subset\Omega$ is called the Pareto Optimal set or efficient solution set. These two sets define the solutions of a MOP, referred to as optimal or efficient solutions. Generally, $PF$ is a surface of dimension $m-1$ in $\mathbb{R}^{m}$.
@@ 2.1 Region of Interest
Figure 1: Preference cone $\mathcal{C}$ with its axis $\mathbf{v}$ and opening angle $\theta$, objective space $F(\Omega)$, Pareto Front ($PF$) and ROI in bi-dimensional space.
@@ 2.2 DWU algorithm
[…]
@@
2:  $N_{\mathcal{X}}\subseteq\mathcal{X}$ such as $F(N_{\mathcal{X}})$ is the set of non-dominated solutions.
@@
[…]
@@ 3 Proposed Approach
$\mathcal{C}\mbox{-}FrontLevel=\begin{cases}\begin{tabular}[]{@{}cc@{}}$FrontLevel+P_{\alpha,\theta}(\phi)$,&$\mbox{if }\phi>\theta$\\ &\\ $FrontLevel$,&$\mbox{if }\phi\leq\theta$\end{tabular}\\ \end{cases}$ (7)
@@
$\mathcal{C}\mbox{-}w_{d}(x,x^{\prime})=\begin{cases}\begin{tabular}[]{@{}c@{}c@{}}$\displaystyle\frac{\|x-x^{\prime}\|}{|d(x)-d(x^{\prime})|+1}-P_{\beta,\theta}(\phi)$,&$\mbox{if }x\mbox{ or }x^{\prime}\notin\mathcal{C}$\\ &\\ $\displaystyle\frac{\|x-x^{\prime}\|}{|d(x)-d(x^{\prime})|+1}$,&$\mbox{if }x\mbox{ and }x^{\prime}\in\mathcal{C}$\end{tabular}\\ \end{cases}$ (8)
@@ 6 Conclusion
The proposed methodology successfully obtains high-quality solutions in the objective space within the predefined Region of Interest with good convergence. It also ensured that the optimal solutions exhibited significantly higher diversity than those obtained by the modified NSGAII algorithm. Defining the Region of Interest(ROI) using the preference cone $\mathcal{C}$ is aviabled alternative to other methods due to its ease of adjusting the direction and extent of the RO, as wellIasd itsstraightforwardy incorporation into existing algorithms.

### v3

@@ 2 Background
$PF$ $=\{\mathbf{y}\in F(\Omega);\nexists~\mathbf{x}\in\Omega\text{ such that }F(\mathbf{x})\prec\mathbf{y}\}$ (2)
$PO$ $=\{\mathbf{x}\in\Omega;F(\mathbf{x})\in PF\}$ (3)
The set $PF\subset F(\Omega)$ is called the Pareto Front, and $PO\subset\Omega$ is called the Pareto Optimal set or efficient solution set. These two sets define the solutions of a MOP, referred to as optimal or efficient solutions. Generally, $PF$ is a surface of dimension m1m-1 in $\mathbb{R}^{m}$.
@@ 2.1 Region of Interest
Figure 1: Preference cone $\mathcal{C}$ with its axis $\mathbf{v}$ and opening angle $\theta$, objective space $F(\Omega)$, Pareto Front (PFPF) and ROI in bi-dimensional space.
@@ 2.2 DWU algorithm
where $\|\cdot\|$ is $L_{2}$ norm-based distance. Finding the Pareto optimal set with maximum dispersion is to solve the following problem
@@
[…]
@@
wherein $s(x)$ in the population $\mathcal{X}$} is defined by the number of solutions in $\mathcal{X}$ it dominates and, $\|\cdot\|$ is $L_{2}$ norm-based distance (Euclidean norm) take up in the decision space.
@@ 3 Proposed Approach
wherein $P_{\alpha,\theta}(\phi)=\left\lfloor{e^{\left(\alpha(\phi-\theta)\right)}}\right\rfloor$ is the penalization function ($\lfloor\star\rfloor$ is the floor function (rounds to the nearest integer less than or equal to)), with $\theta$ (opening angle of the preference cone $\mathcal{C}$) and $\alpha$ (penalty intensity) parameters. $FrontLevel$ refers to the level of non-domination of the solution within the population.
@@
[…]
@@ 6 Conclusion
The proposed methodology successfully obtains high-quality solutions in the objective space within the predefined Region of Interest with good convergence. It also ensured that the optimal solutions exhibited significantly higher diversity than those obtained by the modified NSGAII algorithm. Defining the Region of Interest(ROI) using the preference cone $\mathcal{C}$ is an available alternative to other methods due to its ease of adjusting the direction and extent of the ROI, as well as its straightforward incorporation into existing algorithms.